What is the slope of (1,6) and (7,-2)
Answer by checkley77(12844)

(1,6) and (7,-2)
Slope=(y2-y1)/*x2-x1)
Slope=(-2-6)/(7-1)
Slope=-8/6=-4/3 ans.